|3x+6|=|7x+3| .. what do the x's equal for each side- my teacher showed me but i just cant figure it out
Answer by stanbon(75887)

|3x+6|=|7x+3|
----
Solve 3x+6 = 7x+3 and solve 3x+6 = -(7x+3)
4x = 3 or 3x+6 = -7x-3
x = 3/4 or 10x = -9
---
x = 3/4 or x = -9/10
===============================
Cheers,
Stan H.
=================